The mean annual cost of car insurance is Rs 9450 and standard deviation of Rs 250. What is the probability that the average insurance cost of a sample of 25 cars will be between Rs 9400 and Rs 9500?



1
Expert's answer
2022-01-31T15:27:36-0500

X~N("\mu,\frac{\sigma^2}{n})"

"given~\mu=9450,\sigma = 250,n=25\\
p(\frac{9400-9450}{\frac{250}{\sqrt{25}}}<z<\frac{9500-9450}{\frac{250}{\sqrt{25}}})\\
p(-1<z<1)\\=0.6827"